What is the final balance for the investment? $20,000 for 3 years at 4% compounded annually.

Answer by G8TORS(10)

yr. 1 - 20,000 x .04 = 800 i
yr. 2 - 20,800 x .04 = 832 i
yr. 3 - 21,632 x .04 = 865.28 i
total yeild at the end of year 3 is 22,497.28 or 2,497.28 over investment